1 If you are moving at 50mila and you brake with a=-5m/s , how long does it take to come to a complete stop.



distance do -2) from you the previous problem, what advance before stopping.
Answer #1

Answ091– Given dara PS 4 = 50 milh = 0.0138 mis a=-5 m/s² U V=O we have V-atat 0 = 0.0138+(-5)t t = 0.0138 -0.00275 2) s-ut +